OpenStudy (anonymous):

Angelo is solving this problem. 342 × 405 What are the partial products Angelo will need to find the product? A. 1,710 and 13,680 B. 1,710 and 1,368 C. 1,710 and 136,800 D. 1,710, 3,420, and 136,800

OpenStudy (anonymous):

@Surana

OpenStudy (muscrat123):

it is c

OpenStudy (muscrat123):

do u know how to solve?

OpenStudy (surana):

This isn't really my strong suit. Sorry.

OpenStudy (surana):

@Nnesha ?

OpenStudy (anonymous):

Which number sentences show how to solve the problem? The bowling club donated 4 crates of orange juice for the neighborhood picnic. The soccer club donated 12 crates of mixed juices for the picnic. Each crate holds 32 juice boxes. How many juice boxes were donated altogether for the picnic? Choose all answers that are correct. A. 4 + 12 = 16 and 16 × 32 = 512 B. 4 × 32 = 128 and 128 × 12 = 336 C. 4 × 32 = 128, 12 × 32 = 384, and 128 + 384 = 512 D. 4 + 12 = 16 and 16 × 21 = 336

OpenStudy (anonymous):

@muscrat123

OpenStudy (muscrat123):

yes?

OpenStudy (anonymous):

@muscrat123 can u help me

OpenStudy (muscrat123):

ok so each container holds 32 juice boxes and the bowling club donated 4 crates. each means to multiply. do the same thing for the soccer club donation

OpenStudy (anonymous):

Which operations are the most efficient to use to solve the problem and what is the solution? The community center just built an outdoor lap pool. The new outdoor pool has an area of 432 square feet and a length of 24 feet. The width of the new outdoor pool is 2 feet longer than the width of the center’s indoor pool. What is the width of the community center’s indoor pool? Choose exactly two answers that are correct. A. multiplication and subtraction B. division and subtraction C. 16 ft D. 1,726 ft

OpenStudy (anonymous):

@muscrat123 can u help me